# Multiple Instance Verification

### v1

@@ 4.1.1 Multi-head Cross Attention Pooling
$O^{P}_{j}$ $=MatMul\Bigg{[}\underbrace{as_{j}\Big{(}v^{query}W_{j},v^{target}W_{j}\Big{)}}_{1\times N},\underbrace{LN_{j}\Big{(}v^{target}W_{j}\odot MHSCE_{j}(v^{query})\Big{)}}_{N\times D}\Bigg{]},$ (8a)
$O^{Q}_{j}$ $=\underbrace{LN_{j}\Big{(}v^{query}W_{j}\odot MHSCE_{j}(v^{query})\Big{)}}_{1\times D},\qquad j=1,2,\dots,h$ (8b)
where $h$ is the number of heads, $D=\frac{C}{h}$, $W_{j}\in\mathbb{R}^{C\times D}$ are learnable weights of the multi-head linear projections shared by $\mathbf{Q},\mathbf{K},\mathbf{V}$, and the subscript $j$ of $as_{j}(\cdot,\cdot)$, $MHSCE_{j}(\cdot)$, $LN_{j}(\cdot)$ indicates the $j^{th}$ head. $MatMul$ and $\odot$ denote matrix and element-wise multiplication respectively.
Next, we provide details of the functions $as(\cdot,\cdot)$ and $MHSCE(\cdot)$ for an individual head within the multi-head setting, omitting the subscript $j$ for brevity.
4.1.2 Components of Cross Attention Pooling
VEMA is one of the two attention functions we propose. The rationale is to construct an adaptation of the existing multiplicative attention popularized by the transformer [34]. Additionally, VEMA creates an adjustment using channel-wise variance (across instances within a bag). Intuitively, when all instances in a bag look similar to each other, channels that make them more different, i.e., channels with higher variance, may make it easier to identify key instances. Accordingly, we design the “excitation by variance” of channels, which can be viewed as an additional feature selection (or channel selection) mechanism in attention. We expect it to produce more unevenly-distributed attention scores, and hence more accurate predictions of the key instances, a hypothesis to be tested by the experiments in Section 6. Formally, we have
@@ 4.1.2 Components of Cross Attention Pooling
where $Y$ and $Z$ have shape $(1\times D)$ and $(N\times D)$ respectively, and $\delta\in\mathbb{R}^{1\times D}$. $\delta$ is modelled as a gating mechanism, $\delta=sigmoid\Big{[}relu\Big{(}(variance(v^{target})-1)R\Big{)}S\Big{]}$, where $variance(v^{target})\in\mathbb{R}^{1\times C}$ denotes the channel-wise variance of $v^{target}$, $R\in\mathbb{R}^{C\times C},S\in\mathbb{R}^{C\times D}$ are learnable parameters (with $S$ projecting channels to a single head), and $sigmoid$, $relu$ are element-wise activation functions.
Our motivation for the second attention function DBA is to establish an equally effective but simpler alternative to VEMA. Considering multiplicative attention as a generalized “cosine similarity”, we surveyed alternative similarity measures, which led to DBA as a new attention function. More precisely, the attention scores of DBA are driven by the negative sum of weighted distances between corresponding channels of vectors in $\mathbf{K}$ and $\mathbf{Q}$. In this paper, we use $L_{1}$-distance, noting that other distance metrics are also possible. The DBA attention-score function is
@@
[…]
Finally, to adopt SCE, the same “squeeze” and “excitation” functions are shared in both Equations (8a) and (8b), prior to element-wise multiplications with $(v^{target}W_{j})$ and $(v^{query}W_{j})$ respectively. We apply SCE in a multi-head context, thus naming it “MHSCE”. More precisely, MHSCE’s “excitation” function is the following gating mechanism, in which $mean(x)\in\mathbb{R}^{1\times C}$, denoting the channel-wise mean of $x$, is the “squeeze” function: $MHSCE(x)=sigmoid\Big{[}relu\Big{(}mean(x)J\Big{)}M\Big{]}$, where $J\in\mathbb{R}^{C\times C},M\in\mathbb{R}^{C\times D}$ are learnable parameters, and $sigmoid$, $relu$ are element-wise activation functions. We emphasize that the same $x=v^{query}$, as well as the same learnable parameters, in $MHSCE(x)$ are shared between Equations (8a) and (8b), that is, in both Siamese twins. Also note that $MHSCE(\cdot)$ is multi-headed, because $M$ projects all the channels to a head.
@@ 4.2 Baseline and benchmark models
To provide a baseline for all models, we create a simple model for MI verification. We also adapt SOTA methods from the MIL literature to develop three benchmark models that enable model comparison. The three SOTA methods are: “gated-attention based MIL” (GABMIL) from [17], “pooling by multi-head attention” (PMA) from [21], and “multi-head self attention” (MSA) used in [11, 30]. While the benchmark models are all attention-based, we also experiment with some non-attention-based MI-pooling methods in Section 6, namely MI-Net [36] and bi-direction LSTM [35]. The architectures of the benchmark and attention-based baseline models are depicted in Figure 3 and explained as follows. The architectures of non-attention-based models are similar to that of the benchmark using GABMIL (Figure 3(b))—that is, $v^{P}$, the non-attention-based bag-level representation of $v^{target}$, and $v^{Q}=v^{query}$ are the two outputs of the Siamese twins—and are omitted here for brevity.
@@ 4.2.1 The baseline
Because the baseline model aggregates a bag based on its elements’ scores without an explicit bag-level representation, it is analogous to an extension of instance-level MIL methods, e.g., mi-Net of [36]. Implicitly, it nevertheless still has a bag-level representation for the target bag, which is the feature vector of the instance whose corresponding pair attains the maximum score. More precisely, the baseline can be viewed as a very simple attention model since taking the maximum implements an attention mechanism using the $\mathbbm{1}_{argmax}$ function instead of the $softmax$ function used by other models, where $\mathbbm{1}_{argmax}$ denotes an indicator function providing a vector based on the set of indices returned from the $argmax$ function. [^7] In case there is more than one maximum, the output of $\mathbbm{1}_{argmax}$ is normalized by $L_{1}$-normalization to sum to 1, i.e., $\frac{|x_{n}|}{\sum_{n}|x_{n}|}$. Like in other attention-based methods, the (implicit) bag-level representation in the baseline model is also the attention-score weighted sum of the instances’ representations:
@@
This baseline is limited, due to (1) its simplified architecture with very few learnable parameters, (2) the $argmax$ attention function that allows only those instances that yield maxima to represent the entire bag, and (3) parameters $\alpha$ that are shared by both the Siamese-twin metric and the attention mechanism. Nonetheless, this simple model captures the essence of what we call “cross-attention” because it directly includes the query instance $v^{query}$ as part of attention. Hence, this model is a useful baseline for model comparisons.
@@ 4.2.2 Benchmark using GABMIL
As the first benchmark, we adopt the MI pooling method of Gated-Attention-based MIL (GABMIL) from [17] to aggregate $v^{target}$, which computes one branch of the Siamese twins $v^{P}$ as the attention-weighted sum of $v^{target}$, and another as $v^{Q}=v^{query}$ (cf. Figure 3(b)).
@@ 4.2.3 Benchmark using PMA
The second benchmark applies an MI pooling method based on the transformer decoder, called “Pooling by Multi-head Attention” (PMA) by [21]. PMA has the same architecture as the transformer decoder (without positional encoding and dropout), used to aggregate $v^{target}$.
[21] defines $PMA_{k}$ by applying multi-head attention on a learnable set of $k$ seed vectors $\in\mathbb{R}^{k\times C}$. We initially used $PMA_{1}$ to aggregate $v^{target}$: $v^{P}=PMA_{1}(v^{target};P)$, $v^{Q}=v^{query}$, where $P\in\mathbb{R}^{1\times C}$ is a learnable seed vector. However, exploiting the Siamese-twin structure, we found it useful to treat the query’s feature vector $v^{query}$ as another seed vector when computing $v^{Q}$. More precisely, we stack $v^{query}$ on $P$, and compute the outputs of the Siamese twins as follows (see also Figure 3(c)),
@@
We found that $PMA_{2}$ yields slightly better classification performance than $PMA_{1}$. Therefore, we adopt $PMA_{2}$ as the method for the experiments presented in this paper.
@@ 4.2.4 Benchmark using MSA
Finally, there is an extensive range of applications and variations of transformer encoders in the literature, e.g., [11, 30]. The key component of these is multi-head self-attention (MSA), characterized by the full self-attention of all the instances within a bag to themselves.
MSA per se cannot aggregate multiple instances, because it retains the same positions in the output sequence as in the input sequence. To use MSA as an MI pooling method, the common practice is to include an additional, special token “[CLS]” in the input sequence, whose embedding vector, $v^{[CLS]}$, is learnable. One can treat the vector at the position of “[CLS]” in the output sequence as the aggregation of the rest of the inputs.
Analogous to $PMA_{2}$, we may treat $v^{query}$ as a feature vector of another special token. We then apply the same MSA (without positional encoding and dropout) twice, on both $v^{[CLS]}$ and $v^{query}$. This gives rise to the Siamese twins (in Figure 3(d)) as
@@
Following common practice to stack multiple MSA layers when aggregating $v^{target}$, see, e.g., [30], we stacked two layers of MSA—adding more layers did not appear to add value in our experiments.
@@
We note that none of the three benchmark methods has a natural mechanism to incorporate $v^{query}$ as part of attention when producing the bag-level representation $v^{P}$. Even though the Siamese-twin similarity metric may induce a reference to $v^{query}$ in attention-weighted aggregation of $v^{target}$, this relationship is not modeled explicitly, which may make it less effective for the identification of key instances and for classification. Section 6 provides empirical evidence that supports this hypothesis: the benchmarks are not better, and sometimes significantly worse, than the simple baseline.
5 Evaluating the quality of the explanations
@@
Attention-based MIL models, e.g., [17, 30], reflect the importance of the ability to identify “key instances” using attention scores extracted from models. Following the literature, we also pay attention to the evaluation of explanation quality in the context of MI verification. This is possible when the ground truth of key instances is available in the datasets, as is the case in our study.
@@
Unlike previous literature that considers qualitative demonstrations of explanation quality, this paper focuses on quantitative evaluation. More precisely, for each positive exemplar in the test dataset, we match the ground truth of its key instances’ identifiers to a model’s attention scores. When there are multiple heads, we follow standard practice by averaging the attention scores across all heads to identify the key instances. Figure 4 provides an illustration of such matching for a test QMNIST exemplar. This process is related to evaluation for “weakly-supervised object localization” in CV domains [18, 5], from which we employ two threshold-independent ranking metrics, namely pixel “area under ROC curve” (AUROC), and pixel “average precision” (AP).
@@
Figure 4: An illustration of the quantitative evaluation of explanations. Ground-truth key instances, highlighted by red rectangles, are unavailable during training.
@@
The threshold independence comes from the definitions of AUROC and AP—they are summary statistics of ROC and PR curves (see e.g., [9]), on which each point evaluates the classification performance based on a single decision threshold. Consequently, by summarizing the entire curves, both metrics measure the performance across all possible classification thresholds and thus are threshold-independent. The reasons to require threshold independence in our case are twofold: one, avoid controversies regarding how thresholds are determined to predict key instances (see, for example, [5]); two, for attention models, selecting thresholds may be less straightforward because attention scores must sum to $1$, and the bag size $N$ varies across exemplars. Therefore, sensible thresholds may need to be dynamic for exemplars with different sizes, a subject that is beyond the scope of this paper.
@@
We adopt the same computation as pixel AUROC and pixel AP, except that we replace “pixel” with “instance”. Accordingly, we call them instance AUROC (i-AUROC) and instance AP (i-AP), both of which assess explanations for a single exemplar. To summarize the two metrics for the entire test dataset, we take the average of i-AUROC (average i-AUROC) and i-AP (average i-AP), respectively, across all exemplars in the data.
@@
The example in Figure 4 shows that i-AUROC and i-AP accurately reflect quality of the explanations for an individual exemplar—their values are consistent with the rankings of the attention scores with respect to the ground truth of two key instances. More importantly, in this example, all models correctly predict the exemplar label (=1), but the explanation quality differs. Together with more results in Section 6.2, this demonstrates that even if a model misidentifies key instances, it will quite possibly make a correct classification, and vice versa. Therefore, classification performance alone is insufficient to fully evaluate MI models: it should be complemented by quantitative assessment of explanations, whenever possible.
@@
Note that we omit the models adapted from non-attention-based MIL methods when assessing explanations because they are either unable to identify the key instances [36], or require complicated post-processing to do so [35]. We additionally omit the MSA benchmark when evaluating quality of the explanations, for two reasons. One, there is inconsistency regarding how to properly extract attention scores in the literature relating to the transformer encoder. For instance, [11] used “attention rollout” whereas [30] used the final MSA-layer to obtain attention scores. It is not yet clear which way should be adopted. Two, as MSA is not a natural pooling method, its extracted attention scores include components that are irrelevant for MI pooling due to the self-attention mechanism. It is also unclear how to better post-process (e.g., re-normalize) the extracted attention scores to obtain meaningful explanations.

### v2

@@ 4.1.1 Multi-head Cross-Attention Pooling
$O^{P}_{j}$ $=MatMul\Bigg{[}\underbrace{as_{j}\Big{(}v^{query}W_{j},v^{target}W_{j}\Big{)}}_{1\times N},\underbrace{LN_{j}\Big{(}v^{target}W_{j}\odot MHCE_{j}(v^{query})\Big{)}}_{N\times D}\Bigg{]},$ (8a)
$O^{Q}_{j}$ $=\underbrace{LN_{j}\Big{(}v^{query}W_{j}\odot MHCE_{j}(v^{query})\Big{)}}_{1\times D},\qquad j=1,2,\dots,h$ (8b)
where $h$ is the number of heads, $D=\frac{C}{h}$, $W_{j}\in\mathbb{R}^{C\times D}$ are learnable weights of the multi-head linear projections shared by $\mathbf{Q},\mathbf{K},\mathbf{V}$, and the subscript $j$ of $as_{j}(\cdot,\cdot)$, $MHCE_{j}(\cdot)$, $LN_{j}(\cdot)$ indicates the $j^{th}$ head. $MatMul$ and $\odot$ denote matrix and element-wise multiplication.
Next, we provide details of the functions $as(\cdot,\cdot)$ and $MHCE(\cdot)$ for an individual head within the multi-head setting, omitting the subscript $j$ for brevity.
4.1.2 Components of Cross-Attention Pooling
VEMA is one of the two attention functions proposed in this paper. The rationale is to construct a suitable variant of the “multiplicative attention” function popularized by the transformer (Vaswani et al. (2017)). To this end, VEMA creates an adjustment using channel-wise variance (across instances within a bag). Intuitively, when all instances in a bag look similar to each other, channels that make them more different, i.e., channels with higher variance, may make it easier to identify key instances. Accordingly, we design the “excitation by variance” of channels, which can be viewed as a feature selection (or channel selection) mechanism in attention. We expect it to produce more unevenly distributed attention scores, and hence more accurate predictions of the key instances, a hypothesis to be tested by the experiments in Section 6. Formally, we have
@@ 4.1.2 Components of Cross-Attention Pooling
where $Y$ and $Z$ have shape $(1\times D)$ and $(N\times D)$ respectively, and $\delta\in\mathbb{R}^{1\times D}$. $\delta$ is modeled as a gating mechanism,
$\delta=sigmoid\Big{[}relu\Big{(}(variance(v^{target})-1)R\Big{)}S\Big{]},$ (10)
where $variance(v^{target})\in\mathbb{R}^{1\times C}$ denotes the channel-wise variance of $v^{target}$, $R\in\mathbb{R}^{C\times C}$ and $S\in\mathbb{R}^{C\times D}$ are learnable parameters (with $S$ projecting channels to a single head), and sigmoidsigmoid and relurelu are element-wise activation functions.
Our motivation for the second attention function DBA is to establish an equally effective but simpler alternative to VEMA. Considering multiplicative attention as a generalized “cosine similarity”, we surveyed alternative similarity measures, which led to DBA as a new attention function. More precisely, the attention scores of DBA are driven by the negative sum of weighted distances between corresponding channels of $\mathbf{K}$ and $\mathbf{Q}$. In this paper, we focus on $L_{1}$-distance, although other distance metrics are also possible.
The DBA attention-score function is
@@
where $Y$ and $Z$ are broadcast to the same shape $(1\times N\times D)$, $|\cdot|$ yields element-wise absolute value, $\beta\in\mathbb{R}^{1\times D}$ is a vector of learnable parameters for a single head (with different $\beta$ for different heads), and $c=\sqrt{\frac{4}{\pi}}D$ and $s=\sqrt{(2-\frac{4}{\pi})D}$ are two constant scalars, broadcast as $(1\times N)$ to be compatible with Equation (11). Like the scaling factor in the transformer (Vaswani et al. (2017)), $c$ and $s$ are normalizing factors to maintain numerical stability.777cc, $s$ are the mean and standard deviation of $\sum_{j}^{D}|a_{j}-b_{j}|$ assuming all $a_{j},b_{j}$ are i.i.d. $\sim\mathcal{N}(0,1)$.
To elucidate the rationale behind DBA, we note that distance metrics can be used to form variation measures (or measures of dispersion). To see this, if we average, across all instances in a bag, the channel-wise distances between the bag ($\mathbf{K}$) and the query ($\mathbf{Q}$), we obtain a measure of $\mathbf{K}$’s variation around $\mathbf{Q}$” for each channel. For example, we get “mean absolute deviation around $\mathbf{Q}$” if we use $L_{1}$-distance, and “variance around $\mathbf{Q}$” if we use $L_{2}$-distance. Hence, channels with higher variation around $\mathbf{Q}$ automatically contribute more to the attention score in Equation (11), because high variation implies greater distance. Therefore, DBA is similar in spirit to VEMA but provides a simplification—comparing Equation (11) to (9), we can see DBA contains no explicit formula of variation, even though it implicitly achieves the same effect, and with fewer learnable parameters ($C$ vs. $2C^{2}$). [^8] The actual implementation of VEMA also includes a bias term, resulting in $2(C^{2}+C)$ parameters. Geometrically, distances are also more robust to changes of data origin, than the (non-centered) cosine similarity. The efficacy of this simplification and its robustness, reflected in classification and key instance detection accuracy, is to be examined in Section 6.
Finally, for the MHCE module, the “excitation” function (Hu et al. (2018)) is a gating mechanism akin to the “gated linear unit”(Dauphin et al. (2016)), which is well-known for its ability to enable feature selection. Co-excitation is to apply the same “excitation” function driven by the query to both Siamese twins—that is, to activate simultaneous selection of features for both query and target bag. More concretely, an “excitation” function is shared by both Equations (8a) and (8b), prior to element-wise multiplications with $(v^{target}W_{j})$ and $(v^{query}W_{j})$ respectively. We adopt “co-excitation” in a multi-head context, thus naming it “MHCE”. Formally, MHCE’s “excitation” function is:
$MHCE(x)=sigmoid\Big{[}relu\big{(}xJ\big{)}M\Big{]},$ (12)
where $J\in\mathbb{R}^{C\times C}$ and $M\in\mathbb{R}^{C\times D}$ are learnable parameters, sigmoidsigmoid and relurelu are element-wise activation functions, and $x\in\mathbb{R}^{1\times C}$. We emphasize that the same $x=v^{query}$, as well as the same learnable parameters, in $MHCE(x)$ are shared between Equations (8a) and (8b). Note that $MHCE(\cdot)$ is multi-headed, because $M$ projects all the channels to a head.
@@ 4.2 Baseline and benchmark models
To provide a baseline for all models, we create a simple model for MIV, inspired by Lemma 2 (Section 3.2). We also adapt SOTA methods from the MIL literature to develop three benchmark models that enable model comparison. The three SOTA methods are: “Gated-Attention Based MIL” (GABMIL) from Ilse et al. (2018), “Pooling by Multi-head Attention” (PMA) from Lee et al. (2019), and “Multi-head Self Attention” (MSA) used in Shao et al. (2021). While the benchmark models are all attention-based, we also experiment with some non-attention-based MI-pooling methods in Section 6, namely MI-Net from Wang et al. (2018) and bi-direction LSTM from Wang et al. (2020). The architectures of the baseline and attention-based benchmark models are depicted in Figure 4 and explained in the following. The architectures of non-attention-based models are similar to that of the benchmark using GABMIL (Figure 4(b))—that is, $v^{P}$, the non-attention-based bag-level representation of $v^{target}$, and $v^{Q}=v^{query}$ are the two outputs of the Siamese twins—and are omitted here for brevity.
@@ 4.2.1 The baseline
Because the baseline model aggregates a bag based on its elements’ scores without an explicit bag-level representation, it is analogous to an extension of instance-level MIL methods, e.g., mi-Net of Wang et al. (2018). Implicitly, it nevertheless still has a bag-level representation for the target bag, which is the feature vector of the instance whose corresponding pair attains the maximum score. More precisely, the baseline can be viewed as a very simple attention model, since taking the maximum implements an attention mechanism using the $\mathbbm{1}_{argmax}$ function instead of the softmaxsoftmax function used by other models, where $\mathbbm{1}_{argmax}$ denotes an indicator function providing a vector based on the set of indices returned from the argmaxargmax function. [^9] In case there is more than one maximum, the output of $\mathbbm{1}_{argmax}$ is normalized by $L_{1}$-normalization to sum to 1, i.e., $\frac{|x_{n}|}{\sum_{n}|x_{n}|}$. Like in other attention-based methods, the (implicit) bag-level representation in the baseline model is also the attention-score weighted sum of the instances’ representations:
@@
This baseline model is limited, due to (1) its simplified architecture with very few learnable parameters, (2) the argmaxargmax attention function that allows only those instances that yield maxima to represent the entire bag, and (3) parameters $\alpha$ that are shared by both the Siamese-twin metric and the attention mechanism. Nonetheless, this simple model explicitly handles bags of paired instances and captures the essence of a representation of MIV data in Lemma 2 (Section 3.2). Hence, this model is a useful baseline for model comparison.
@@ 4.2.2 Benchmark using GABMIL
As the first benchmark, we adopt the MI pooling method of Gated-Attention-Based MIL (GABMIL) from Ilse et al. (2018) to aggregate $v^{target}$, which computes one of the Siamese twins $v^{P}$ as the attention-weighted sum of $v^{target}$, and another as $v^{Q}=v^{query}$ (Figure 4(b)).
@@ 4.2.3 Benchmark using PMA
The second benchmark applies an MI pooling method based on the transformer decoder, called “Pooling by Multi-head Attention” (PMA) by Lee et al. (2019). PMA has the same architecture as the transformer decoder (without positional encoding and dropout), used to aggregate $v^{target}$. Lee et al. (2019) define $PMA_{k}$ by applying the transformer’s multi-head cross-attention on a learnable set of $k$ seed vectors $\in\mathbb{R}^{k\times C}$.
We initially used $PMA_{1}$ to aggregate $v^{target}$: $v^{P}=PMA_{1}(v^{target};P)$, $v^{Q}=v^{query}$, where $P\in\mathbb{R}^{1\times C}$ is a learnable seed vector. However, exploiting the Siamese-twin structure, we found it useful to compute $v^{Q}$ as an aggregation of $v^{target}$ by treating $v^{query}$ as a seed vector. More precisely, we stack $v^{query}$ on $P$, and compute the outputs of the Siamese twins as follows (see also Figure 4(c)):
@@
We found that $PMA_{2}$ yields slightly better classification performance than $PMA_{1}$. Hence, we adopt $PMA_{2}$ as the benchmark in the experiments presented in our study.
Notably, $PMA_{2}$ consists of paired PMAs that adapt to Siamese networks to solve a verification problem, where PMA is constructed using the transformer’s cross-attention mechanism. In this sense, this benchmark can be seen as an adaptation of standard, “transformer-style” cross attention—without fundamentally changing its core architecture—to the MIV setting.
@@ 4.2.4 Benchmark using MSA
Finally, there is an extensive range of applications and variations of transformer encoders in the literature (e.g., Shao et al. (2021); Dosovitskiy et al. (2021)). The key component of these methods is multi-head self-attention (MSA), characterized by the full self-attention of all the instances within a bag to themselves.
MSA per se cannot aggregate multiple instances because it retains the same positions in the output sequence as in the input sequence. To use MSA as an MI pooling method, the common practice is to include an additional, special token, “[CLS]”, in the input sequence, whose embedding vector, $v^{[CLS]}$, is learnable. One can treat the vector at the position of “[CLS]” in the output sequence as the aggregation of the rest of the inputs.
Analogously to $PMA_{2}$, we treat $v^{query}$ as a feature vector of another special token. We then apply the same MSA (without positional encoding and dropout) independently to both $v^{[CLS]}$ and $v^{query}$. This gives rise to the Siamese twins (in Figure 4(d)) as
@@
Following common practice to stack multiple MSA layers when aggregating $v^{target}$ (see, e.g., Shao et al. (2021)), we stacked two layers of MSA—adding more layers did not appear to add value—in our experiments.
@@
We highlight that none of the three benchmark methods has a natural mechanism to incorporate $v^{query}$ as part of attention when producing the bag-level representation $v^{P}$. Although the Siamese-twin similarity metric may induce a reference to $v^{query}$ in the attention-weighted aggregation of $v^{target}$, this relationship is not modeled explicitly, which may make it less effective for the identification of key instances and for classification. Section 6 provides empirical evidence that supports this hypothesis: the benchmarks are not better, and sometimes significantly worse, than even the simple baseline.
5 Evaluating performance of key instance detection
@@
Attention-based MIL models, for example, the models presented in (Ilse et al. (2018); Shao et al. (2021)), reflect the importance of the ability to identify “key instances” using attention scores extracted from models. Following this work, and similar other studies (including Jia et al. (2021); Early et al. (2022); Choe et al. (2023)) we quantitatively evaluate the ability to identify key instances using attention scores, in the context of MIV. This is possible when the ground truth of key instances is available in the datasets, as is the case in our study.
@@
More precisely, for each positive exemplar in the test dataset, we match the ground truth of its key instances’ identifiers to a model’s attention scores. When there are multiple heads, we follow standard practices by averaging the attention scores across all heads to identify the key instances. Figure 5 provides an illustration of such matching for a test QMNIST exemplar. This process is related to evaluation for “weakly-supervised object localization” in CV domains (Jia et al. (2021); Choe et al. (2023)), from which we employ two threshold-independent ranking metrics, namely pixel “area under ROC curve” (AUROC), and pixel “average precision” (AP).
@@
Figure 5: An illustration of the quantitative evaluation of key instance detection. Note that ground-truth key instances, highlighted by red rectangles, are unavailable during training.
@@
The threshold independence comes from the definitions of AUROC and AP—they are summary statistics of ROC and PR curves (see, e.g., Davis and Goadrich (2006)), on which each point evaluates the classification performance based on a single decision threshold. Consequently, by summarizing entire curves, both metrics measure the performance across all possible classification thresholds and thus are threshold-independent. The reasons to require threshold independence in our case are twofold: one, to avoid controversies regarding how thresholds are determined to predict key instances (see, for example, Choe et al. (2023)); two, for attention-based models, selecting thresholds may be less straightforward because attention scores must sum to $1$, and the bag size $N$ varies across exemplars. Therefore, sensible thresholds may need to be dynamic for exemplars with varying sizes, a subject that is beyond the scope of this paper.
@@
We adopt the same computation as pixel AUROC and pixel AP, except that “pixel” becomes “instance” in our case. Accordingly, we call them instance AUROC (i-AUROC) and instance AP (i-AP), both of which assess key instance detection performance for a single exemplar. To summarize the two metrics for the entire test dataset, we take the average of i-AUROC (average i-AUROC) and i-AP (average i-AP), respectively, across all exemplars in the data.
@@
The example in Figure 5 shows that i-AUROC and i-AP accurately reflect the performance when detecting key instances for an individual exemplar—their values are consistent with the rankings of the attention scores with respect to the ground truth of two key instances. More importantly, in this example, all models correctly predict the exemplar label (=1), but prediction of key instances differs. Together with more results in Section 6.2, this demonstrates that even if a model misidentifies key instances, it will quite possibly make a correct classification, and vice versa. Therefore, classification performance alone is insufficient to assess MI models: it should be complemented by quantitative evaluation of key instance detection performance, whenever possible.
@@
Note that we omit the models using non-attention-based MIL methods when assessing key instance detection because they are either unable to identify the key instances (Wang et al. (2018)), or require complicated post-processing to do so (Wang et al. (2020)). We additionally omit the MSA benchmark when considering key instance detection, for two reasons. One, there is inconsistency regarding how to properly extract attention scores in the literature relating to the transformer encoder. For instance, Dosovitskiy et al. (2021) used “attention rollout” whereas Shao et al. (2021) used the final MSA-layer to obtain attention scores. It is not yet clear which way should be adopted. Two, as MSA is not a natural pooling method, its extracted attention scores include components that are irrelevant for MI pooling due to the self-attention mechanism. It is also unclear how to better post-process (e.g., re-normalize) the extracted attention scores to properly identify key instances.